# Reasoning Language Models: A Blueprint

### v1

@@ A.2 Monte Carlo Tree Search (MCTS)
Edge statistics The MCTS algorithms stores the following three values for every edge $s,a$ in the search tree:
@@
  • $r(s,a)=r(s,a,s^{\prime})$ the reward received after taking the action $a$ in the state $s$ leading to state $s^{\prime}$.
@@ B.3.2 Value Model (VM)
V-Value Model (V-VM). One variant of a value model is the V-Value Models which predicts the expected cumulative future reward of a state, denoted as $V(s)$. This is equivalent to the state value function in reinforcement learning, which evaluates the long-term potential of the current state $s$. A key advantage of V-VMs is their global perspective, as they aggregate future rewards across all possible trajectories originating from the current state. However, V-VMs do not explicitly evaluate individual actions, which may limit their utility in step-level decision-making. Additionally, V-values are often ill-defined at terminal states, where rewards may substitute for state values during training.
@@ C.1.1 Setup and Notation
We now introduce the necessary notation to understand MCTS and the training pipeline algorithmically. Moreover, we will propose our design choices for x1.
MDP Design x1 defines a MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},p,r,\gamma)$ for RLMs as defined in Appendix A.1. We chose to adopt a sparse reward structure, where only transitions into terminal states yield non-zero rewards. In these terminal states, the final reasoning step provides an answer that can be extracted and, during training, compared to a given solution (commonly referred to as the golden answer). Therefore, the reward objective is tied to the correctness of the final answer. In many reasoning tasks like math, access is typically limited to the final answer, with no ground truth available for intermediate reasoning steps (which are often also non-unique). As a result, the reward structure is inherently sparse. We assign a reward of $r(s_{T},a_{T},s_{T+1})=1$ for correct answers and $r(s_{T},a_{T},s_{T+1})=-1$ for incorrect ones where $T+1$ denotes the terminal state. All other MDP components are chosen as in Appendix A.1. $\gamma$ can be chosen freely but we propose to select values between $[0.95,1]$ to not over-penalize long reasoning sequences. Since rewards are sparse and take values between -1 and 1, the state value function simplifies to
$V_{\pi_{\theta}}(s_{t})=\mathbb{E}_{\pi}\left[\gamma^{T-t}r(s_{T},a_{T})\mid s_{t}\right]\quad\in[-1,1]$ (4)
and the state action function breaks down to:
@@
$Q_{\pi}(s,a)\stackrel{{\scriptstyle}}{{=}}\begin{cases}r(s_{T},a_{T}),&\text{if }t=T\\ \gamma V_{\pi}(s_{t+1}),&\text{otherwise}\end{cases}\quad\in[-1,1]$ (5)
Due to our definition of rewards, we obtain bounded state value and state action functions – a great property that we will use in the MCTS design and the design of the value model architecture.
@@
MCTS Design We define the MCTS tree as in Appendix A.2 as $T=(N,E)$, where $N$ is a set of nodes, and $E$ is the set of edges. Furthermore, each state $s_{i}$ defines a reasoning sequence: $s_{i}=(z_{0},\dots,z_{i})$ which starts from an initial question $z_{0}$ followed by reasoning steps $z_{1},z_{2},...,z_{i}$. The MCTS search process constructs the tree with the aim of achieving a terminal node $s_{T+1}=(z_{0},...,z_{T+1})$ that contains a sound sequence of reasoning steps to a final answer $z_{T+1}$. Individual reasoning steps $z_{j}$ are obtained by querying the policy LM with the context $s_{j-1}$. Thus MCTS enables an efficient search over the space of reasoning sequences, towards a valid final answer. To simplify, we use the notation of a node-edge-node relationship denoted by $(s,a^{\prime},s^{\prime})$, where $s$ represents the origin node, $a^{\prime}$ describes the action corresponding to an edge, and $s^{\prime}$ denotes the target node. This notation symbolically ties the action and the target state together, as the action uniquely identifies the target state and is therefore indicative of it.
Figure 13: Example MCTS generated tree of reasoning sequences.
The policy model Given a question $z_{0}$, the algorithm uses a parametric policy model $\pi_{\theta}$ (a pretrained LM) to generate candidate reasoning steps. In our implementation, special tokens are used to delimit individual reasoning steps. To help the policy model distinguish between individual reasoning steps, a special token called the ’End of Intermediate Step’ (eois) is defined to mark the end of a reasoning step. In addition, we need a method to detect if a state is terminal for the MCTS algorithm. By introducing a separate token for the end of intermediate steps, we can leverage the end-of-sequence (eos) token from the RLM itself. The eos token naturally signals the conclusion of a model’s reasoning process, indicating that the generated response is complete. The presence of an eos token can then be used to identify terminal states.
@@ C.1.2 MCTS Algorithm
We now present the algorithmic steps of a Monte Carlo Tree Search variant similar to AlphaZero as implemented in the x1 reasoning framework. The MCTS search operates in two distinct modes. The core difference is that, during training, a final answer verifier evaluates and scores terminal nodes, providing a true reward signal that is backpropagated through the tree. This reward serves as a reliable learning signal for the value model $Q_{\varphi}$. During inference, however, the verifier is unavailable, and decisions rely solely on the value model.
Notation. We chose to store all values in nodes instead of edges, which defines the following set of statistics saved for each node $s^{\prime}$:
  • $N(s^{\prime})$ the visit count of node $s^{\prime}$ by the algorithm,
  • $q(s^{\prime})$ the running estimate of the q-value for the state $s^{\prime}$,
  • $\mathbb{T}(s^{\prime})$ the binary value indicating if the node $s^{\prime}$ is terminal (00: not terminal, $1$: terminal).
Note that the q-value mathematically correspond to the value $Q(s,a^{\prime})$ where $s$ is the parent of $s^{\prime}$ and $s^{\prime}$ is the result of taking action $a^{\prime}$ from state $s$. This corresponds to traversing the directed edge $(s,a^{\prime},s^{\prime})$, as the q-value describes the quality of taking action $a^{\prime}$ in state $s$ to reach $s^{\prime}$. Since only $s^{\prime}$ can uniquely characterize this transition, we chose it as the location to store the value. Additionally, we denote the set of children of a state $s$ as $\mathcal{C}(s)$ and we denote the children as $s_{c}$. The transition from $s$ to its child $s_{c}$ is then tied to the action $a_{c}$. Inversely, we denote the parent of a node $s$ as $s_{p}$ and the corresponding action that connects $s$ and $s_{p}$ as $a_{p}$.
@@
Expansion. We opted for an explorative expansion phase, where $M$ nodes are appended to the selected leaf. In more greedy frameworks, only the highest-ranked candidate is appended to the MCTS. However, one of the major challenges when searching with RLMs is the lack of diversity and the extensive exploration required to cover the search space effectively. We found that an explorative expansion, where suboptimal nodes are also appended, provides additional diversity, as such nodes can sometimes lead to successful directions needed to solve complex tasks. To achieve this, each selected node is prompted $M$ times to the policy model to generate potential reasoning steps. Additionally, we employ diverse beam search to further enhance exploration across nodes, leveraging diversity in the decoding strategy. Once generated, all nodes are scored using the Q-value model $q(s_{c})=Q_{\varphi}(s,a_{c}),\forall s_{c}\in\mathcal{C}(s)$ and corresponding actions $a_{c}$. Finally, terminality is assessed by checking whether the newly generated children include the (eos) token, which signals a final answer.
Backpropagation. The backpropagation step serves to propagate information from the terminal nodes back to their ancestors. In our implementation, the running estimates of the Q-values for the nodes along the selection path are updated using a convex combination of their previous values and the discounted child values, weighted by an empirical distribution derived from the visit counts. For a node $s$, this update rule is defined as:
$q_{\text{new}}(s)=$ $(1-\alpha)q_{\text{old}}(s)+\alpha\left(\sum_{s_{c}\mathcal{C}(s)}\pi_{\text{MCTS}}(a_{c}\mid s)\gamma q(s_{c})\right),$
where we look at the node-edge-node tuples $(s,a_{c},s_{c})$ and $s_{c}\in\mathcal{C}(s)$. The ”MCTS” distribution $\pi_{\text{MCTS}}$ is defined over the visit scores of nodes as follows:
$\pi_{\text{MCTS}}(a_{c}\mid s)=\frac{N(s_{c})}{\sum_{s_{\tilde{c}}\in\mathcal{C}(s)}N(s_{\tilde{c}})}.$
To motivate this formula, note that the visit scores of $s_{c}$ are directly tied to the action $a_{c}$, due to the deterministic transitions in the MDP. Consequently, the probability of visiting the edge corresponding to action $a_{c}$ in state $s$ is determined by the visit scores of the resulting state $s_{c}$, compared to the visit scores of all other child states $s_{\tilde{c}}$ accessible from $s$. This update rule effectively corresponds to a SARSA update for sparse rewards, utilizing $M$ values weighted by the visit scores. Additionally, the visit counts are incremented for each node along the path during this process.
True Reward Propagation. We propose to improve the quality of the q-values by propagating the real final rewards back through the tree when a terminal state is reached. During training, terminal nodes can be evaluated against a reference golden answer $g^{*}$ using an external verifier. A reasoning sequence leading to a valid final answer can then be evaluated with a better evaluator than the q-value model. We suggest to replace the initialization of the running estimate of a terminal state with the true reward based on the evaluation of the external verifier which in terminal states is equal to the q-value see Eq. 5. The reward is then backpropagated via the q-values through the tree with our backpropagation operator. This adjustment anchors the q-value model predictions with real reward signals and prevents the q-value model predictions to diverge. Due to the need for access to a true answer, this correction can not be applied at inference. Therefore, we do not use this adjustment step in the expansion phase of the inference mode.
Best Path Selection. After $N$ iterations, MCTS will have formed a tree of explored reasoning trajectories. The path with the highest terminal value estimate is ultimately selected and its final reasoning step is returned as the final solution.
@@
Input: Policy model $\pi_{\theta}$, value model $Q_{\varphi}$, question $z_{0}$, golden answer $g^{*}$, binary correctness verifier $\Gamma$, number of MCTS iterations $N$, number of children expanded in every selection phase $M$, exploration constants $c_{1},c_{2}$, MCTS reward-value convexity constant $\alpha$.
Output: Search tree $\mathcal{T}=(\mathcal{N},\mathcal{E})$ with the best path $\tau^{*}$.
1:  $s\leftarrow s_{0}\leftarrow(z_{0})$Initialize root node}
@@
10:        $s_{c}\leftarrow\operatorname*{arg\,max}_{s_{c}\in\mathcal{C}(s)}q(s_{c})+\frac{\sqrt{N(s)-1}}{1+N(s_{c})}$
11:        $\qquad\qquad\qquad\qquad\qquad\qquad\quad\cdot\left(c_{1}+\log{\frac{N(s)+c_{2}}{c_{2}}}\right)$ )
@@
16:        $z_{c}\leftarrow(t^{1},\ldots t^{z_{c}})\sim\pi_{\theta}${Sample a new reasoning step}
@@
18:        $q(s_{c})\leftarrow Q_{\varphi}(s,a_{c})$Predict with the Q-VM}
@@
20:        if $s_{c}$ terminal then
21:           $\mathbb{T}(s_{c})\leftarrow 1$Mark as terminal}
22:           $r(s,a_{c})\leftarrow\begin{cases}1,&\text{if }\Gamma(s_{c},g^{*})=1,\\ -1,&\text{if }\Gamma(s_{c},g^{*})=0.\end{cases}$Check for correctness with golden answer}
23:           $q(s_{c})\leftarrow r(s,a_{c})$Overwrite by true reward}
24:        end if
25:        $\mathcal{N}\leftarrow\mathcal{N}\cup\{s_{c}\}$Add the node to the tree}
26:        $\mathcal{E}\leftarrow\mathcal{E}\cup\{(s,s_{c})\}$Add the edge to the tree}
27:     end for
28:     ————– Backpropagation ——————————–
29:     while $s\neq s_{0}$ do
30:        $N(s)\leftarrow N(s)+1$Update the visit count}
31:        $q(s)\leftarrow(1-\alpha)q(s)+\alpha\gamma\sum_{s_{c}\in\mathcal{C}(s)}\pi_{MCTS}(a_{c}|s)q(s_{c})$
32:        {Update the value}
33:        $s\leftarrow s_{p}$Move to the parent}
34:     end while
35:     $i\leftarrow i+1$
36:  end while
37:  Best Path Selection:
38:  Select the best reasoning sequence $s^{*}_{T}$.
39:
40:  return  $s^{*}_{T}$, all reasoning sequences $\{s^{(i)}_{j}\}_{j}$
@@
C.2 Training Pipeline
To adequately employ the MCTS-based reasoning scheme introduced in the Appendix C.1, the policy model must be fine-tuned to generate responses in the format of semantically-relevant reasoning steps. The value model – a q-value model in our case – must be trained to accurately estimate the values of the sequences of reasoning steps.
@@ C.2 Training Pipeline
C.3 Phase 1: Supervised Fine-Tuning
The first phase focuses on preparing the policy and value models to generate and evaluate reasoning trajectories effectively. This is achieved by supervised fine-tuning (SFT) training on a dataset of example sequences of reasoning steps (where intermediate reasoning steps are terminated by an ”End of Intermediate Step” eois token). The objective is twofold: (1) to fine-tune the policy model $\pi_{\theta}$ to produce semantically coherent reasoning steps, and (2) to train the q-value model $Q_{\varphi}$ to accurately assign scalar scores to reasoning trajectories, distinguishing between high-quality and suboptimal reasoning paths.
@@ C.3 Phase 1: Supervised Fine-Tuning
C.3.1 Datasets Generation and Preparation
Dataset for SFT of the Policy. Performing SFT of the policy requires a dataset of high-quality reasoning sequences denoted as $D_{\text{SFT}}=\{\left(x_{\text{SFT}}^{(i)},\,y_{\text{SFT}}^{(i)}\right)\}$. Each pair in the dataset consists of a prompt $x_{\text{SFT}}^{(i)}$ composed of a sequence of reasoning steps (for example $x_{\text{SFT}}^{(i)}=(z_{0},...,z_{j})$), and a target completion $y_{\text{SFT}}^{(i)}=z_{j+1}$ which is the subsequent reasoning step or final answer. Appendix D contains a detailed account of the dataset creation and processing. It covers how the special eois token is appended to reasoning steps mark the end of a step during inference.

### v2

@@ A.2 Monte Carlo Tree Search (MCTS)
Edge statistics The MCTS algorithm stores the following three values for every edge $s,a$ in the search tree:
@@
  • $r(s,a)=r(s,a,s^{\prime})$ the reward received after taking the action $a$ in the state $s$ leading to the state $s^{\prime}$,
  • $\beta(s,a)$ the terminality function indicating if the action $a$ leads to a terminal state.
@@ B.3.2 Value Model (VM)
V-Value Model (V-VM). One variant of a value model is the v-value model which predicts the expected cumulative future reward of a state, denoted as $V(s)$. This is equivalent to the state value function in reinforcement learning, which evaluates the long-term potential of the current state $s$. A key advantage of V-VMs is their global perspective, as they aggregate future rewards across all possible trajectories originating from the current state. However, V-VMs do not explicitly evaluate individual actions, which may limit their utility in step-level decision-making. Additionally, v-values are often ill-defined at terminal states, where rewards may substitute for state values during training.
@@ C.1.1 Setup and Notation
We will now present the details of the training pipeline of x1.
MDP Design x1 assumes the MDP following the definition presented in Appendix A.1 with the $\gamma$ values between $[0.95,1]$ to avoid over-penalizing long reasoning sequences. In the RLM setup, the state space and action space of the underlying MDP constitute a tree in which every state $s$ other than the starting state $s_{0}$ has exactly one action $a_{s}$ leading to it. This allows us to simplify the notation by omitting actions wherever it’s clear from the context that we are referring to only action leading to a given. For every action $a$ leading from the state $s$ to the state $s^{\prime}$ we will write:
$\pi(s^{\prime}\mid s):=\pi(a_{s^{\prime}}|s)$
$r(s^{\prime}):=r(s,a,s^{\prime})$
$q(s^{\prime}):=q(s,a)$.
$\tau:=(s_{0},s_{1},\dots,s_{T+1})$
The final reasoning step in the terminal state contains the RLM’s answer to the original query. The final answer is compared to the ground truth solution, commonly referred to as the golden answer. This matches the common setup in many reasoning tasks and math problems, where no ground truth and no reward source is available for the intermediate reasoning steps.
Consider a trajectory $\tau:=(s_{0},s_{1},\dots,s_{T+1})$. We assign a reward of $r(s_{T+1})=1$ if the last reasoning step in the final state $s_{T+1}$ contains the correct answer and $r(s_{T+1})=-1$ otherwise. The state value function simplifies to
$V_{\pi}(s_{t})$ $=\mathbb{E}_{\pi}\left[\gamma^{T-t}r(s_{T+1})\right]\quad\in[-1,1]$ (4)
and the state action function can be rewritten as:
@@
$Q_{\pi}(s_{t})\stackrel{{\scriptstyle}}{{=}}\begin{cases}r(s_{T+1}),&\text{if }t=T+1\\ \gamma V_{\pi}(s_{t+1}),&\text{otherwise}\end{cases}\quad\in[-1,1]$ (5)
hence both the value and the state-action value functions are bounded between -1 and 1 for all states and state-action pairs.
@@
MCTS Design We define the MCTS tree as in Appendix A.2 as $\mathcal{T}=(N,E)$, where $N$ is a set of nodes, and $E$ is the set of edges. We use the notation of a node-edge-node relationship denoted by $(s,a^{\prime},s^{\prime})$ where $s$ represents the origin node, $a^{\prime}$ describes the action corresponding to an edge, and $s^{\prime}$ denotes the target node. This notation symbolically ties the action and the target state together, as the action uniquely identifies the target state and is therefore indicative of it.
The policy model We use a pretrained LM with parameters $\theta$ as a policy model and denote it $\pi_{\theta}$. The model autoregressively generates a sequence of tokens. We use a special token ’End of Intermediate Step’ (eois) to indicate the end of the reasoning step. We use a standard end-of-sequence (eos) token to indicate the end of the final reasoning step concluding the reasoning trajectory.
@@ C.1.2 MCTS Algorithm
We now present the algorithmic steps of a Monte Carlo Tree Search variant similar to AlphaZero as implemented in the x1 reasoning framework. The MCTS search operates in two distinct modes: training and inference. The core difference is that, during training, a final answer verifier evaluates and scores the final reasoning steps, providing a true reward signal that is backpropagated through the MCTS tree. This reward serves as a reliable learning signal for the value model $Q_{\varphi}$. During inference, however, the verifier is unavailable, and decisions rely solely on the value model.
Notation. We chose to store all values in nodes instead of edges, which defines the following set of statistics saved for each node $s$:
  • $N(s)$ the visit count of node $s$
  • $q(s)$ the running estimate of the q-value of the transition leading to state $s$,
  • $\beta(s)$ the binary terminality function, returns $1$ if the node $s$ is terminal 00 otherwise.
@@
Expansion. We append $M$ nodes to the selected leaf, $M$ being a hyperparameter. One of the major challenges in applying RLMs is maintaining the diversity of reasoning paths. By adding $M$ nodes, we increase the exploration of alternative reasoning paths.
Backpropagation. The backpropagation step serves to propagate information from the terminal nodes back to their ancestors. In our implementation, we update the running estimates of the q-values using the following formula:
$q(s)\leftarrow$ $(1-\alpha)q(s)+\alpha\gamma\left(\sum_{s_{c}\mathcal{C}(s)}w_{s}(s_{c})\cdot q(s_{c})\right),$
where we look at the node-edge-node tuples $(s,a_{c},s_{c})$ and $s_{c}\in\mathcal{C}(s)$. The weights $w_{s}(s_{c})$ for combining the children q-values are defined over the visit scores of the nodes as follows:
$w_{s}(s_{c})=\frac{N(s_{c})}{\sum_{s_{\tilde{c}}\in\mathcal{C}(s)}N(s_{\tilde{c}})}.$
True Reward Propagation. We improve the quality of the q-values by propagating the real final rewards back through the tree when a terminal state $s_{T+1}$ is reached. During training, terminal nodes can be evaluated against a reference golden answer $g^{*}$ using an external verifier. For actions leading to terminal states, the associated reward is equal to the q-value see Eq. 5. Therefore, instead of using the prediction of the q-value model, we initialize $q(s_{T+1})$ with the true reward $r(s_{T+1})$ based on the evaluation of the external verifier. The reward is then backpropagated via the q-values through the tree with our backpropagation operator. This adjustment anchors the q-value model predictions with real reward signals and prevents the q-value model predictions to diverge.
Best Path Selection. After $N$ iterations, MCTS will have formed a tree in which every path corresponds to one of the explored reasoning trajectories. The final reasoning step in a path with the highest terminal value estimate is returned as the final solution.
@@
Input: Policy model $\pi_{\theta}$, value model $Q_{\varphi}$, question $z_{0}$, golden answer $g^{*}$, binary correctness verifier $\Gamma$, number of MCTS iterations $N$, number of children expanded in every selection phase $M$, exploration constants $c_{1},c_{2}$, Backpropagation weight $\alpha$.
Output: Search tree $\mathcal{T}=(\mathcal{N},\mathcal{E})$ containing the best path $\tau^{*}$.
1:  $s_{0}\leftarrow(z_{0})$Initialize root node}
@@
10:        {Select child $s_{c}\in\mathcal{C}(s)$ with highest selection score}
11:        $s_{c}\leftarrow\operatorname*{arg\,max}\limits_{s_{c}\in\mathcal{C}(s)}q(s_{c})+\frac{\sqrt{N(s)-1}}{1+N(s_{c})}\left(c_{1}+\log{\frac{N(s)+c_{2}}{c_{2}}}\right)$ )
@@
16:        $z_{c}\leftarrow(t^{1},\ldots t^{M_{z_{c}}})\sim\pi_{\theta}${Sample a new reasoning step}
@@
18:        $q(s_{c})\leftarrow Q_{\varphi}(s)$Predict with the Q-VM}
@@
20:        $\beta(s_{c})\leftarrow 0$Initialize terminality function}
21:        if $s_{c}$ terminal then
22:           $\beta(s_{c})\leftarrow 1$Mark as terminal}
23:           $r(s_{c})\leftarrow\begin{cases}1,&\text{if }\Gamma(s_{c},g^{*})=1,\\ -1,&\text{if }\Gamma(s_{c},g^{*})=0.\end{cases}$Check for correctness to determine the reward}
24:           $q(s_{c})\leftarrow r(s_{c})$Overwrite by true reward}
25:        end if
26:        $\mathcal{N}\leftarrow\mathcal{N}\cup\{s_{c}\}$Add the node to the tree}
27:        $\mathcal{E}\leftarrow\mathcal{E}\cup\{(s,s_{c})\}$Add the edge to the tree}
28:     end for
29:     ————– Backpropagation ——————————–
30:     while $s\neq s_{0}$ do
31:        $N(s)\leftarrow N(s)+1$Update the visit count}
32:        $q(s)\leftarrow(1-\alpha)q(s)+\alpha\gamma\sum_{s_{c}\in\mathcal{C}(s)}w_{s}(s_{c})q(s_{c})$
33:        {Update the value}
34:        $s\leftarrow s_{p}$Move to the parent}
35:     end while
36:     $i\leftarrow i+1$
37:  end while
38:  Best Path Selection:
39:  Select the best reasoning sequence $s^{*}_{T}$.
40:
41:  return  $s^{*}_{T}$, all reasoning sequences $\{s^{(i)}_{j}\}_{j}$
@@
C.2 Training Phase 1
Overall Training Pipeline. To adequately employ the MCTS-based reasoning scheme introduced in the Appendix C.1, the policy model must be fine-tuned to generate responses in the format of semantically-relevant reasoning steps. The value model – a q-value model in our case – must be trained to accurately estimate the values of the sequences of reasoning steps.
@@ C.2 Training Phase 1
Phase 1: Supervised Fine-Tuning. The first phase focuses on preparing the policy and value models to generate and evaluate reasoning trajectories effectively. This is achieved by supervised fine-tuning (SFT) training on a dataset of example sequences of reasoning steps (where intermediate reasoning steps are terminated by an ”End of Intermediate Step” eois token). The objective is twofold: (1) to fine-tune the policy model $\pi_{\theta}$ to produce semantically coherent reasoning steps, and (2) to train the q-value model $Q_{\varphi}$ to accurately assign scalar scores to reasoning trajectories, distinguishing between high-quality and suboptimal reasoning paths.
@@
C.2.1 Datasets Generation and Preparation
Dataset for SFT of the Policy. Performing SFT of the policy requires a dataset of high-quality reasoning sequences denoted as $D_{\text{SFT}}=\{\left(x_{\text{SFT}}^{(i)},\,y_{\text{SFT}}^{(i)}\right)\}$. Each pair in the dataset consists of a prompt $x_{\text{SFT}}^{(i)}$ composed of a sequence of reasoning steps (for example $x_{\text{SFT}}^{(i)}=(z_{0}^{(i)},...,z_{j}^{(i)})$), and a target completion $y_{\text{SFT}}^{(i)}=z_{j+1}^{(i)}$ which is the subsequent reasoning step or final answer. Appendix D contains a detailed account of the dataset creation and processing. It covers how the special eois token is appended to reasoning steps mark the end of a step during inference.